OpenStudy (anonymous):

what is the equivalent fractions for 3/4

OpenStudy (heisenberg):

\[\frac{3}{4} * \frac{n}{n} , n \in \mathbb{Z}\]

OpenStudy (heisenberg):

that last part means, where n is an integer (a non zero integer in this case). there are an infinite number of equivalent fractions